40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 29 18:24:11 2011, jamie, Update, SUS, ITMY stuck to OSEMs? 
    Reply  Tue Aug 30 14:44:06 2011, kiwamu, Update, SUS, ITMY released without opening chambers 
       Reply  Fri Sep 2 17:57:18 2011, kiwamu, Update, SUS, Re: ITMY released without opening chambers 
Message ID: 5326     Entry time: Tue Aug 30 14:44:06 2011     In reply to: 5320     Reply to this: 5338
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ITMY released without opening chambers 

The ITMY mirror was released. The OSEM readouts became healthy.

 

To see what is going on, I changed the PIT DC bias slider on ITMY from 0.8 to -1 or so, and then the optic started showing a free swinging behavior.

If there were no responses to the DC bias, I was going to let people to open the chamber to look at it closer, but fortunately it released the optic.

Then I brought the slider back to 0.8, and it looked still free swinging. Possibly the optic had been stacked on some of the OSEMS as Jamie expected.

Quote from #5320

ITMY, which is supposed to be fully free-swinging at the moment, is displaying the tell-tale signs of  being stuck to one of it's OSEMs. 

Do we have a procedure for remotely getting it unstuck?  If not, we need to open up ITMYC and unstick it before we pump.

 

ELOG V3.1.3-